Ford Tops manufactures hats for baseball teams. Ford has fixed costs of $323,000 per quarter and sells each hat for $38. If the variable cost per hat is $19.00, how many hats must Ford sell each quarter to break even? (Do not round intermediate calculations.)
Answer #1

Contribution margin=Sales-Variable cost

=38-19=$19 per hat

Breakeven=Fixed cost/Contribution margin

=323,000/19

=17000 hats.